OpenStudy (nirmalnema):

trigonometric equations:- (1-tan@)(1+sin2@)=1+tan@ @=theta

jhonyy9 (jhonyy9):

sin2 @ is sin squared or just sin2 ?

OpenStudy (nirmalnema):

simply sin2@...not squared... =2sin@cos@

jhonyy9 (jhonyy9):

do you know it sure that tan x = (sin x)/cos x yes ?

OpenStudy (nirmalnema):

yup..

jhonyy9 (jhonyy9):

try rewriting this equation subtituting tan x = (sin x)/cos x because you need getting an equation with just one the same unknowed term with theta for example sin theta or cos theta ok?

OpenStudy (nirmalnema):

(cos@-sin@ / cos@)(sin2@)=(cos@+sin@ / cos@)

jhonyy9 (jhonyy9):

yes nice but sin2@ rewrite it how you have wrote in your first answer

OpenStudy (nirmalnema):

(cos@-sin@ / cos@)(1+2sin@cos@)=(cos@+sin@ / cos@)

jhonyy9 (jhonyy9):

how you have got this 1 from 1+2sinxcosx ?

jhonyy9 (jhonyy9):

first you have wrote sin2x = 2sinxcosx

OpenStudy (nirmalnema):

its already in the ques...i forgot to write it

jhonyy9 (jhonyy9):

ok than rewrite it please the correct exercise

OpenStudy (nirmalnema):

its is there 1+sin2@

jhonyy9 (jhonyy9):

yes i see it

jhonyy9 (jhonyy9):

so if you check it again you see that there are two fractions with the same denominator cos x

jhonyy9 (jhonyy9):

cos@ sorry

jhonyy9 (jhonyy9):

yes ?

jhonyy9 (jhonyy9):

(cos@-sin@ / cos@)(1+2sin@cos@)=(cos@+sin@ / cos@) (cos@ -sin@) (cos@ +sin@) ----------- * (1 +2sin@cos@) = ------------ cos@ cos@ so now do you understand it ?

jhonyy9 (jhonyy9):

so and hope you know again sure that sin^2 @ +cos^2 @ =1 yes ?

jhonyy9 (jhonyy9):

than on the left side inside paranthese there is an 1 what you can rewrite it like sin^2 @ +cos^2 @ ok ?

jhonyy9 (jhonyy9):

so and in this case you will get the formula a^2 +2ab +b^2 what you know that is equal (a+b)^2

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

so than will get (cos@ -sin@)(cos@ +sin@)^2 =(cos@ +sin@) divide both sides by (cos@ +sin@) and will get (cos@ -sin@)(cos@ +sin@)=1 but you see that on the left side there is formula (a+b)(a-b)=a^2 -b^2